OpenStudy (anonymous):

Robert earned $4 each hour doing chores for his neighbors. He also earned $20 working for his mom.. Altogether Robert earned $80. write an equation that shows the situation..use your equation to find the number of hours Robert worked doing chores for his neighbors... please help

OpenStudy (anonymous):

@BBoyd01 Medal? Fan?

OpenStudy (anonymous):

can u help me with this problem

OpenStudy (anonymous):

4*20=80 doesn't it? Now find the hours using the question

OpenStudy (anonymous):

60/4+20=80

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

4+20 is 24 not 80 but multiply and get 80 not 24!!

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Your welcome @BBoyd01

OpenStudy (anonymous):

15 hours because 60/4=15 then plus the 20 = 80

OpenStudy (anonymous):

@Brandon1016 can u come and check ur question! :D Than u can help ur friend ! :) http://openstudy.com/study#/updates/5306963ce4b022471a65a135